Circular No. 8373

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004dd IN NGC 124 Further to IAUC 8372, J. Graham and W. Li report the LOSS discovery

of an apparent supernova (mag ∼ 17.1) on an unfiltered KAIT image taken on July 12.44 UT. The new object was confirmed on an earlier image taken on July 1.45 (mag ∼ 17.0). SN 2004dd is located at α = 0^{h}27^{m}52^{s}.98, δ = −1^{o}48′46′′.3 (equinox 2000.0), which is 9′′.9 east and 9′′.5 south of the nucleus of NGC 124. A KAIT image taken on 2003 Dec. 28.25 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2004de IN NGC 3054 L. A. G. Monard reports his discovery of an apparent supernova on

unfiltered CCD images taken on July 15.703 (at mag ∼ 14.5 ± 0.2) and 17.688 UT (mag ∼ 14.6 ± 0.1) with a 0.30-m reflector near Pretoria, South Africa. The new object is located at α = 9^{h}54^{m}27^{s}.48, δ = −25^{o}41′34′′.7 (equinox 2000.0), which is 13′′ west and 37′′ north of the nucleus of NGC 3054. Nothing is visible at this position on images taken by Monard on June 20.7 (limiting red mag 17.7) or on his earlier images, and nothing is visible on a red Digitized Sky Survey image (limiting mag 20).

RINGS OF URANUS I. de Pater, University of California at Berkeley; H. B. Hammel, Space

Science Institute; and S. Gibbard, Lawrence Livermore National Laboratory, report the detection of a faint sheet of material in Uranus’ ring system inside (and clearly displaced from) rings 4, 5, and 6. This sheet of material is visible between ∼ 35000 and ∼ 41000 km from the center of the planet, peaking in intensity at 39000 km. It must be the same ring as 1986 U 2R (Stone and Miner 1986, Science 233, 39), discovered by Voyager in 1986 at visible wavelengths and large phase angles. It is now for the first time visible from the ground because the rings are slowly ‘closing up’, so that the line-of-sight traverses more ring material. The observations were conducted on 2004 July 3, 4, 8, and 9 UT with the 10-m Keck II Telescope (+ NIRC2/AO) on Mauna Kea, at K′-band (2.2 μm). This sheet of material is visible in single 2-5-min exposures; all data (∼ 1 hr) were combined to derive the specifics given above. The ring opening angle was ∼ 11^{o}, with phase angle 2^{o}.25. As we approach ring plane crossing (in 2007), it is expected that more faint dust sheets in the ring system will become visible.

2004 July 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green